Sam recieved $1500 he gave his daugter 1/3 after that he put 2/5 in savings. He donated the rest how much did he donate?
Answer by ikleyn(52754)

.
he donated  1500 - (1/3)*1500 - (2/5)*1500 = 1500 - 500 - 600 = 400 dollars.